OpenStudy (anonymous):

f(a+2)=square1-(a+2) Help please

OpenStudy (anonymous):

what does square1 mean?

OpenStudy (anonymous):

Sorry didn't know how put square sign in. \[f(a+2)=\sqrt{1-(a+2)}\]

OpenStudy (anonymous):

so what do you need done here?

OpenStudy (anonymous):

Can I simplify?

OpenStudy (anonymous):

yes, f(a+2) = sqrt(1-a-2) = sqrt(-1-a) that's about it...

OpenStudy (anonymous):

Thank you